A heavier mass m1 and a lighter mass m2 are 18.5 cm apart and experience a gravitational force of attraction that is 9.00 10-9 N in magnitude. The two masses have a combined value of 5.45 kg. Determine the value of each individual mass.
Answer #1

here,

the separation between the masses , r = 18.5 cm = 0.185 m

m1 + m2 = 5.45 ......(1)

and

the gravitational force , F = G * m1 * m2 /r^2

9.00 * 10^-9 = 6.67 * 10^-11 * m1 * m2 /0.185^2 ......(2)

from(1) and (2)

m1 = 4.40 kg

m2 = 1.05 kg

the heavier mass is 4.40 kg

the lighter mass is 1.05 kg